# Boosting CLIP Adaptation for Image Quality Assessment via Meta-Prompt Learning and Gradient Regularization

### v1

@@ 3.3 Meta-Prompt Pre-training Module
where $\bm{\beta}$ represents the outer learning rate. Meta-learning effectively trains the learnable prompts $\bm{\theta=[{\theta_{T}},{\theta_{V}}]}$, ensuring their generalization across various image distortions.
@@
HyperIQA (Su et al. 2020) 0.648 0.725 0.790 0.615 0.710 0.776 0.790 0.824 0.909 0.892 0.912 0.929 0.102 0.302 0.379
MetaIQA (Zhu et al. 2020) 0.604 0.626 0.669 0.618 0.620 0.660 0.784 0.849 0.894 0.840 0.880 0.919 0.332 0.348 0.371
DEIQT (Qin et al. 2023) 0.667 0.718 0.812 0.638 0.682 0.754 0.821 0.891 0.941 0.920 0.942 0.955 0.396 0.410 0.436
MANIQA (Yang et al. 2022) 0.642 0.769 0.797 0.652 0.755 0.810 0.794 0.847 0.874 0.909 0.928 0.957 0.136 0.361 0.470
CONTRIQUE (Madhusudana et al. 2022) 0.695 0.729 0.761 0.733 0.794 0.821 0.840 0.926 0.940 0.891 0.922 0.943 0.379 0.437 0.488
Re-IQA (Saha, Mishra, and Bovik 2023) 0.591 0.621 0.701 0.685 0.723 0.754 0.893 0.907 0.923 0.884 0.894 0.929 0.280 0.350 0.431
CLIP (Radford et al. 2021) 0.664 0.721 0.733 0.736 0.770 0.782 0.841 0.892 0.941 0.896 0.923 0.941 0.254 0.303 0.368
CLIP-IQA (Wang, Chan, and Loy 2023) 0.695 0.738 0.746 0.692 0.743 0.762 - - - - - - - - -
LIQE (Zhang et al. 2023) 0.691 0.769 0.810 0.759 0.801 0.832 0.838 0.891 0.924 0.904 0.934 0.948 - - -
GRepQ (Srinath et al. 2024) 0.760 0.791 0.822 0.812 0.836 0.855 0.878 0.914 0.941 0.926 0.937 0.953 0.390 0.450 0.498
@@
Table 1: SRCC performance comparison of our method with other IQA methods trained on limited labels. Bold indicates the best results, underlined marks the second-best, and the fifth-to-last through second-to-last lines show the CLIP-based IQA.
3.4 Quality-Aware Gradient Regularization
We denote the original semantic CLIP model as $\bm{V^{sem}}$, and the CLIP model obtained from the Meta-Prompt Pre-training Module as $\bm{V^{qua}}$. Recent studies (Zhao et al. 2023) have shown that IQA tasks can be misaligned with the high-level semantic representations of upstream tasks, leading to overfitting and reduced generalization. To mitigate this issue, we introduce Quality-aware Gradient Regularization (QGR), which uses directional guidance from the original CLIP model’s semantic predictions $\bm{V^{sem}}$ to refine quality gradient updates during fine-tuning on IQA tasks. This approach aims to reduce the focus on semantics and enhance the model’s ability to evaluate image quality effectively.
Specifically, to derive a general semantic direction $\bm{G_{sem}}$, we first design a hard prompt following (Ying et al. 2020): ${w_{i}}=$ “a photo of a [class]” representing one of nine categories: “animal”, “cityscape”, “human”, “indoor”, “landscape”, “night”, “plant”, “still-life”, “others”. This prompt ${w_{i}}$ is input into the text encoders of $\bm{V^{sem}}$ and $\bm{V^{qua}}$, generating text features that align with image features from their respective visual encoders. This results in zero-shot semantic prediction probabilities $\bm{p^{sem}(w_{i}|x)}$ and $\bm{p^{qua}(w_{i}|x)}$. To avoid overemphasis on semantics in $\bm{V^{qua}}$, we measure the semantic attention by calculating the KL divergence between $\bm{p^{sem}(w_{i}|x)}$ and $\bm{p^{qua}(w_{i}|x)}$:
$\mathcal{L}_{kl}(V^{qua})=-\sum_{i}p^{sem}(w_{i}|x)\log\frac{p^{qua}(w_{i}|x)}{p^{sem}(w_{i}|x)}.$ (7)
Conversely, we obtain the quality optimization direction by calculating the cross-entropy ${\mathcal{L}_{ce}(V^{qua})}$ between the prediction $\bm{p(q_{\text{high}}|x)}$ and the ground truth $y$, as shown in Eq. 3. We then balance the learning between quality and semantics by adjusting the gradients of the tasks within a shared representation space. The gradients $\nabla\mathcal{L}_{ce}(V^{qua})$ and $\nabla\mathcal{L}_{kl}(V^{qua})$ are represented as $\bm{G_{qua}}$ and $\bm{G_{sem}}$, and the relationship between them can be described in two aspects:
(1) If the angle between $\bm{G_{qua}}$ and $\bm{G_{sem}}$ is less than 90 degrees, it indicates that the optimization directions for IQA quality knowledge and general semantic knowledge are consistent. In such cases, we clip the $\bm{G_{qua}}$ along its component $\bm{G_{\parallel}}$ that parallel to the semantic direction $\bm{G_{{sem}}}$, to modulate the model’s original quality-aware optimization path, thereby preventing overfitting to semantic correlations.
(2) Conversely, if the angle between two areas of knowledge is more than 90 degrees, it means they are heading in different directions during improvement. In such situations, we don’t adjust the quality gradient $\bm{G_{qua}}$, allowing the model to learn quality-aware features according to the original optimization direction. In conclusion, our gradient regularization strategy is mathematically formulated as:
$G_{qgr}=\begin{cases}G_{qua}&\text{if }G_{qua}\cdot G_{sem}\leq 0,\\ G_{qua}-\lambda\frac{G_{qua}\cdot G_{sem}}{\|G_{sem}\|^{2}}G_{sem}&\text{otherwise},\end{cases}$ (8)
Here, $\lambda$ is introduced to generalize the formulation, providing flexibility in controlling the influence of general knowledge. Specifically, $\lambda=1$ projects $\bm{G_{qua}}$ onto the orthogonal direction of $\bm{G_{sem}}$, while $\lambda=0$ reduces QGR to CoOp.
@@ 4.1 Datasets and Evaluation Protocols
We conduct experiments on seven typical BIQA datasets and one AI-generated dataset, AGIQA-3K (Li et al. 2023a) (results in the supplementary materials). The seven typical datasets include LIVEC (Ghadiyaram and Bovik 2015), KonIQ (Hosu et al. 2020), LIVEFB (Ying et al. 2020), and SPAQ (Fang et al. 2020), which feature authentic distortions, and PIPAL (Jinjin et al. 2020), LIVE (Sheikh, Sabir, and Bovik 2006), and CSIQ (Larson and Chandler 2010), which feature synthetic distortions. LIVEC contains 1,162 mobile device images, SPAQ includes 11,125 photos from 66 smartphones, KonIQ has 10,073 images from open sources, and LIVEFB is the largest real-world dataset with 39,810 images. For synthetic distortions, LIVE and CSIQ contain 779 and 866 images with 5 and 6 types of distortions, respectively. PIPAL, a challenging dataset, includes 23,200 images with 40 types of distortions, including GAN-generated artifacts. We use Spearman’s Rank Correlation Coefficient (SRCC) and the Pearson Linear Correlation Coefficient (PLCC) outcomes as metrics to quantify the monotonousness and accuracy of predictions.
LIVE CSIQ LIVEC KonIQ LIVEFB SPAQ
Method PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC
Training Ratio 20% 20% 20% 20% 20% 20% 30% 30% 10% 10% 20% 20%
DEIQT  (Qin et al. 2023) 0.968 0.965 0.885 0.862 0.822 0.792 0.922 0.903 0.624 0.538 0.912 0.908
GRMP-IQA (Ours) 0.972 0.970 0.958 0.951 0.897 0.875 0.936 0.925 0.686 0.604 0.925 0.920
Training Ratio 80% 80% 80% 80% 80% 80% 80% 80% 80% 80% 80% 80%
MetaIQA  (Zhu et al. 2020) 0.959 0.960 0.908 0.899 0.802 0.835 0.856 0.887 0.507 0.540 - -
CONTRIQUE  (Madhusudana et al. 2022) 0.961 0.960 0.955 0.942 0.857 0.845 0.906 0.894 0.641 0.580 0.919 0.914
DEIQT  (Qin et al. 2023) 0.982 0.980 0.963 0.946 0.894 0.875 0.934 0.921 0.663 0.571 0.923 0.919
Re-IQA  (Saha, Mishra, and Bovik 2023) 0.971 0.970 0.960 0.947 0.854 0.840 0.923 0.914 0.733 0.645 0.925 0.918
LIQE  (Zhang et al. 2023) 0.951 0.970 0.939 0.936 0.910 0.904 0.908 0.919 - - - -
CLIP-IQA+  (Wang, Chan, and Loy 2023) - - - - 0.832 0.805 0.909 0.895 0.593 0.575 0.866 0.864
QFM-IQM (Li et al. 2024) 0.983 0.981 0.965 0.954 0.913 0.891 0.936 0.922 0.667 0.567 0.924 0.920
LoDa  (Xu et al. 2024) 0.979 0.975 - - 0.899 0.876 0.944 0.932 0.679 0.578 0.928 0.925
@@ 4.2 Implementation Details and Setups
We build our model on CLIP-B/16 (Radford et al. 2021). During pre-training, we optimize the visual-text prompt, while in fine-tuning, we adjust the last four blocks of the image and text encoders. For competing models, we use either public implementations or re-train them on our datasets. We use 80% of each dataset for training and 20% for testing. To ensure content independence in datasets with synthetic distortions, training and testing sets are split by reference images. For more details, please refer to the Appendix.
Meta Pre-training Stage. We pre-train on TID2013 (Ponomarenko et al. 2015) and KADID-10K (Lin, Hosu, and Saupe 2019), which contain extensive distortion information. We set the learning rates $\alpha$ and $\beta$ to 1e-4 and 1e-2, and train for 50 epochs using Adam (Kingma and Ba 2014).
Fully Supervised Setting. For fully supervised training, we randomly crop each input image into 10 patches of $224\times 224$ resolution and train the model for 9 epochs using AdamW (Loshchilov and Hutter 2019). The learning rate is ${5\times 10^{-6}}$, with a scheduler over 9 decay epochs. The batch size is 16 for LIVEC and 128 for KonIQ.
Few-Shot Learning Setting. In the few-shot setting, we train GRMP-IQA with 50, 100, or 200 samples from 80% of the training images, following the CoOp (Zhou et al. 2022b) schedule. The $\lambda$ parameter is set to 5.
4.3 Few-Shot Performance Comparison
The GRMP-IQA model acquires a substantial amount of image assessment knowledge, allowing it to furnish powerful and rich priors for various IQA scenarios. Consequently, our model can achieve excellent evaluation performance with a relatively small amount of data. Specifically, we randomly select subsets of 50, 100, and 200 samples from the training set for training and evaluate the same test data as the full data supervised learning. We report the median performance obtained across 10 splits. Tab. 1 presents the comparison results in this few-shot setting. It can be observed that our method outperforms the second-best model GRepQ (Srinath et al. 2024) by a significant margin, even though GRepQ is specifically designed for few-shot learning. Furthermore, our approach significantly outperforms MetaIQA, a meta-learning-based method, highlighting the superiority of meta-prompts for rapid adaptation of CLIP to various IQA scenarios. These results clearly demonstrate the strong capability of our method to learn image quality even with limited labels. For comparison, we also present results in the fully-supervised setting in Table 2 as an extreme case.
@@ 4.3 Few-Shot Performance Comparison
4.4 Performance comparison with SOTA
The results of the comparison between GRMP-IQA and other BIQA methods, which include self-training BIQA methods like CONTRIQUE (Madhusudana et al. 2022) and Re-IQA (Saha, Mishra, and Bovik 2023), as well as CLIP-based BIQA methods such as LIQE (Zhang et al. 2023) and CLIP-IQA+ (Wang, Chan, and Loy 2023), are presented in Tab. 2. It can be observed on six of the eight datasets that GRMP-IQA outperforms all other methods in terms of performance. Achieving leading performance on all of these datasets is a challenging task due to the wide range of image content and distortion types. Therefore, these observations confirm the effectiveness and superiority of GRMP-IQA in accurately characterizing image quality. Notably, our method achieved results competitive with SOTA methods by utilizing just 20% of the training data from the training set.
@@ 4.5 Generalization Capability Validation
To assess the generalization capacity of GRMP-IQA, we conducted cross-dataset validation experiments. In these experiments, the model was trained on one dataset (after meta-learning) and subsequently tested on others, without any adjustment of parameters. The outcomes of these experiments are displayed in Tab. 3, which shows the SRCC achieved across five different datasets. Notably, GRMP-IQA outperforms state-of-the-art (SOTA) models in all six experiments involving cross-authentic datasets. This includes significant performance enhancements in the LIVEC and KonIQ datasets. Moreover, GRMP-IQA exhibits strong competitiveness on synthetic datasets such as LIVE and CSIQ.
@@
Table 4: Ablation experiments with Meta-Prompt Pre-training component under zero-shot on three datasets. The best performances are highlighted in boldface.
Component LIVEC KonIQ
Pre-training Meta-learning PLCC SRCC PLCC SRCC
0.825 0.796 0.788 0.764
0.823 0.788 0.792 0.761
0.858 0.828 0.844 0.811
Table 5: Ablation study on meta-learning effectiveness, showing fine-tuning results on LIVEC with 50 samples.

### v2

@@ 3.3 Meta-Prompt Pre-training Module
where $\beta$ represents the outer learning rate. Meta-learning effectively trains the learnable prompts $\bm{\theta}=\textbf{[}\bm{\theta_{T}},\bm{\theta_{V}}\textbf{]}$, ensuring their generalization across various image distortions.
3.4 Quality-Aware Gradient Regularization
We denote the original semantic CLIP model as $\bm{V}^{\text{sem}}$ and the model obtained from the Meta-Prompt Pre-training Module, as $\bm{V}^{\text{qua}}$. The gradient for the IQA task is the quality gradient $\bm{G}_{\text{qua}}$, while the gradient for the semantic direction is the semantic gradient $\bm{G}_{\text{sem}}$. Recent studies [53, 1] have shown that IQA tasks can be misaligned with the high-level semantic representations of upstream tasks, leading to overfitting and reduced generalization. To prevent semantic overfitting, our Quality-aware Gradient Regularization (QGR) follows three key steps: derive a semantic direction $\bm{G}_{\text{sem}}$, compute the quality optimization direction $\bm{G}_{\text{qua}}$, and balance gradients based on their directional relationship.
Specifically, to derive a general semantic direction $\bm{G}_{\text{sem}}$, we first design a hard prompt following [49, 52]: $\bm{w}_{i}=\text{``a photo of a [class]''}$, where [class] represents one of nine categories: “animal”, “cityscape”, “human”, “indoor”, “landscape”, “night”, “plant”, “still-life”, “others”. This prompt $\bm{w}_{i}$ is input into the text encoders of $\bm{V}^{\text{sem}}$ and $\bm{V}^{\text{qua}}$, generating text features that align with image features $\bm{f}^{qua}$ from their respective visual encoders. As a result, we obtain zero-shot semantic prediction probabilities $\bm{p}^{\text{sem}}(\bm{w}_{i}|\bm{x})$ and $\bm{p}^{\text{qua}}(\bm{w}_{i}|\bm{x})$. To measure the semantic alignment between $\bm{V}^{\text{qua}}$ and $\bm{V}^{\text{sem}}$, we compute the KL divergence:
$\mathcal{L}_{\text{kl}}(\bm{V}^{\text{qua}})=-\sum_{i}\bm{p}^{\text{sem}}(\bm{w}_{i}|\bm{x})\log\frac{\bm{p}^{\text{qua}}(\bm{w}_{i}|\bm{x})}{\bm{p}^{\text{sem}}(\bm{w}_{i}|\bm{x})}.$ (7)
The gradient of $\mathcal{L}_{\text{kl}}(\bm{V}^{\text{qua}})$, denoted as $\bm{G}_{\text{sem}}$, represents the general semantic optimization direction. Similarly, the quality direction $\bm{G}_{\text{qua}}$ is obtained by computing the gradient of the cross-entropy loss $\mathcal{L}_{\text{ce}}(\bm{V}^{\text{qua}})$, which compares the predicted probability $\bm{p}(\bm{q}_{\text{high}}|\bm{x})$ with the ground truth $y$, as shown in Eq. 3. Finally, to balance the gradients between quality $\bm{G}_{\text{qua}}$ and semantics $\bm{G}_{\text{sem}}$ within a shared representation space, we analyze their relationship under two cases:
(1) If the angle between $\bm{G}_{\text{qua}}$ and $\bm{G}_{\text{sem}}$ is less than 90 degrees, it indicates that the optimization directions for IQA quality knowledge and general semantic knowledge are consistent. In such cases, we clip the $\bm{G}_{\text{qua}}$ along its component $\bm{G_{\parallel}}$ that parallel to the semantic direction $\bm{G}_{\text{sem}}$, to modulate the model’s original quality optimization path, preventing overfitting to semantic correlations.
(2) Conversely, if the angle between two areas of knowledge is more than 90 degrees, it means they are heading in different directions during improvement. In such situations, we don’t adjust the quality gradient $\bm{G}_{\text{qua}}$, allowing the model to learn quality-aware features according to the original optimization direction. In conclusion, our gradient regularization strategy is mathematically formulated as:
$\bm{G}_{\text{qgr}}=\begin{cases}\bm{G}_{\text{qua}}&\text{if }\bm{G}_{\text{qua}}\cdot\bm{G}_{\text{sem}}\leq 0,\\ \bm{G}_{\text{qua}}-\lambda\frac{\bm{G}_{\text{qua}}\cdot\bm{G}_{\text{sem}}}{\|\bm{G}_{\text{sem}}\|^{2}}\bm{G}_{\text{sem}}&\text{otherwise},\end{cases}$ (8)
Here, $\lambda$ is introduced to generalize the formulation, providing flexibility in controlling the influence of general knowledge. Specifically, $\lambda=1$ projects $\bm{G}_{\text{qua}}$ onto the orthogonal direction of $\bm{G}_{\text{sem}}$, while $\lambda=0$ reduces QGR to CoOp [56].
@@ 3.4 Quality-Aware Gradient Regularization
HyperIQA [41] 0.648 0.725 0.790 0.615 0.710 0.776 0.790 0.824 0.909 0.892 0.912 0.929 0.102 0.302 0.379
MetaIQA [58] 0.604 0.626 0.669 0.618 0.620 0.660 0.784 0.849 0.894 0.840 0.880 0.919 0.332 0.348 0.371
DEIQT [33] 0.667 0.718 0.812 0.638 0.682 0.754 0.821 0.891 0.941 0.920 0.942 0.955 0.396 0.410 0.436
MANIQA [48] 0.642 0.769 0.797 0.652 0.755 0.810 0.794 0.847 0.874 0.909 0.928 0.957 0.136 0.361 0.470
CONTRIQUE [30] 0.695 0.729 0.761 0.733 0.794 0.821 0.840 0.926 0.940 0.891 0.922 0.943 0.379 0.437 0.488
Re-IQA [37] 0.591 0.621 0.701 0.685 0.723 0.754 0.893 0.907 0.923 0.884 0.894 0.929 0.280 0.350 0.431
CLIP w/ Linear Probe [36] 0.664 0.721 0.733 0.736 0.770 0.782 0.841 0.892 0.941 0.896 0.923 0.941 0.254 0.303 0.368
[…]
LIQE [52] 0.691 0.769 0.810 0.759 0.801 0.832 0.838 0.891 0.924 0.904 0.934 0.948 - - -
GRepQ [40] 0.760 0.791 0.822 0.812 0.836 0.855 0.878 0.914 0.941 0.926 0.937 0.953 $0.390^{\dagger}$ $\underline{0.450}^{\dagger}$ $\underline{0.498}^{\dagger}$
@@
Table 1: SRCC performance comparison of our method with other IQA methods trained on limited labels. Bold indicates the best results, underlined marks the second-best, and the fifth-to-last through second-to-last lines show the CLIP-based IQA. $\dagger$ is our reproduction.
LIVE CSIQ LIVEC KonIQ LIVEFB SPAQ
Method PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC
Training Ratio 20% 20% 20% 20% 20% 20% 20% 20% 10% 10% 20% 20%
DEIQT [33] 0.968 0.965 0.885 0.862 0.822 0.792 0.908 0.888 0.624 0.538 0.912 0.908
LoDa [46] - - - - 0.854 0.815 0.923 0.907 - - - -
GRMP-IQA (Ours) 0.972 0.970 0.958 0.951 0.897 0.875 0.931 0.915 0.686 0.604 0.925 0.920
Training Ratio 80% 80% 80% 80% 80% 80% 80% 80% 80% 80% 80% 80%
MetaIQA [58] 0.959 0.960 0.908 0.899 0.802 0.835 0.856 0.887 0.507 0.540 - -
CONTRIQUE [30] 0.961 0.960 0.955 0.942 0.857 0.845 0.906 0.894 0.641 0.580 0.919 0.914
DEIQT [33] 0.982 0.980 0.963 0.946 0.894 0.875 0.934 0.921 0.663 0.571 0.923 0.919
Re-IQA [37] 0.971 0.970 0.960 0.947 0.854 0.840 0.923 0.914 0.733 0.645 0.925 0.918
LIQE [52] 0.951 0.970 0.939 0.936 0.910 0.904 0.908 0.919 - - - -
CLIP-IQA+ [44] - - - - 0.832 0.805 0.909 0.895 0.593 0.575 0.866 0.864
CDINet [54] 0.975 0.977 0.960 0.952 0.880 0.865 0.928 0.916 - - 0.922 0.919
QFM-IQM [22] 0.983 0.981 0.965 0.954 0.913 0.891 0.936 0.922 0.667 0.567 0.924 0.920
LoDa [46] 0.979 0.975 - - 0.899 0.876 0.944 0.932 0.679 0.578 0.928 0.925
@@ 4.1 Datasets and Evaluation Protocols
We conduct experiments on multiple BIQA datasets include LIVEC [7], KonIQ [10], LIVEFB [49], and SPAQ [3], which feature authentic distortions, and PIPAL [12], LIVE [38], and CSIQ [17], which feature synthetic distortions. LIVEC contains 1,162 mobile device images, SPAQ includes 11,125 photos from 66 smartphones, KonIQ has 10,073 images from open sources, and LIVEFB is the largest real-world dataset with 39,810 images. For synthetic distortions, LIVE and CSIQ contain 779 and 866 images with 5 and 6 types of distortions. PIPAL, a challenging dataset, includes 23,200 images with 40 types of distortions, including GAN-generated artifacts. We use Spearman’s Rank Correlation Coefficient (SRCC) and the Pearson Linear Correlation Coefficient (PLCC) as metrics to quantify the monotonousness and accuracy of predictions.
@@ 4.2 Implementation Details and Setups
We build our model on CLIP-B/16 [36]. During pre-training, we optimize the visual-text prompt, while in fine-tuning, only the last four blocks of the image and text encoders are trained. For competing models, we use public results or re-train them under our setup. Each dataset is split into 80%/20% for training and testing, with splits based on reference images to ensure content independence in datasets with synthetic distortions. To ensure robustness, we report the median performance across ten random splits. All experiments are run on four NVIDIA RTX3090 GPUs.
Meta-Prompt Pre-training. We pre-train on TID2013 [32] and KADID-10K [25] dataset, which contain extensive distortion information. We set the learning rates $\alpha$ and $\beta$ to 1e-4 and 1e-2, and train for 50 epochs using Adam [16].
Fully Supervised Learning Setting. We randomly crop each input image into 10 patches of $224\times 224$ resolution and train the model for 9 epochs using AdamW [26]. The learning rate is $5\times 10^{-6}$}, with a scheduler over 9 decay epochs. The batch size is 16 for LIVEC and 128 for KonIQ.
Few-Shot Learning Setting. In the few-shot setting, we follow the approach in a previous study [40] to train our GRMP-IQA model using randomly selected subsets of 50, 100, and 200 samples from the training set. The regularization weight $\lambda$ is fixed at 5, and the training hyperparameters are set according to the schedule proposed by CoOp [56].
4.3 Performance Comparison with SOTA
Our method effectively acquires extensive image quality assessment knowledge, enabling it to provide powerful priors for various IQA scenarios. Tab. 1 and Tab. 2 summarize the comparative results for different experimental settings.
Few-Shot Setting. Benefiting from the quality priors acquired during pre-training, our model achieves superior performance even with limited training data. As shown in Tab. 1, our GRMP-IQA significantly outperforms the second-best model, GRepQ [40], which is specifically designed for few-shot learning. Additionally, our approach exhibits clear advantages over MetaIQA, a meta-learning-based method, underscoring the effectiveness of meta-prompts for rapidly adapting CLIP to diverse IQA tasks. These results validate GRMP-IQA’s strong capability in learning quality-aware feature with limited labeled samples.
Fully Supervised Setting. Tab. 2 compares GRMP-IQA with other BIQA methods under full supervised setting, including self-training approaches like CONTRIQUE [30] and Re-IQA [37], as well as CLIP-based methods such as LIQE [52] and CLIP-IQA+ [44]. GRMP-IQA outperforms almost all competitors on six datasets. Achieving leading performance across these datasets is particularly challenging due to the diverse range of image content and distortion types. Notably, the proposed method achieves competitive results with the state-of-the-art (SOTA) method on several datasets while utilizing only 20% of the training data compared to the fully supervised setting.
@@ 4.3 Performance Comparison with SOTA
Component LIVEC KonIQ
Pre-training Meta-learning PLCC SRCC PLCC SRCC
0.825 0.796 0.788 0.764
0.823 0.788 0.792 0.761
0.858 0.828 0.844 0.811
Table 4: Ablation study on the effectiveness of meta-learning when trained with only 50 samples.
@@ 4.4 Generalization Capability Validation
To evaluate the generalization capacity of GRMP-IQA, we performed cross-dataset validation experiments, where the model was trained on one dataset and tested on others without parameter adjustments. Tab. 3 reports the SRCC results across five datasets. GRMP-IQA consistently outperforms state-of-the-art models in most cross-authentic scenarios, achieving notable improvements on LIVEC and KonIQ dataset. Additionally, it demonstrates strong competitiveness on synthetic datasets like LIVE and CSIQ.
@@
Table 5: Ablation experiments with Meta-Prompt Pre-training component. The best performances are highlighted in boldface.